### v1

# Reasoning or Memorization? Unreliable Results of Reinforcement Learning Due to Data Contamination

Abstract
The reasoning capabilities of large language models (LLMs) have been a longstanding focus of research within the community. Recent works have further enhanced these capabilities by reinforcement learning (RL), and numerous novel methods have been proposed, claiming to achieve significant improvements with minimal or even no external supervision. Surprisingly, some studies even suggest that random or incorrect reward signals can enhance reasoning performance. However, these breakthroughs are predominantly reported on the Qwen2.5 model family and evaluated on well-known mathematical benchmarks such as MATH-500, AMC, and AIME, while often failing to yield comparable gains on other model families such as Llama, which warrants more in-depth investigation. In this work, our empirical analysis shows that, despite the Qwen 2.5 series attaining superior mathematical-reasoning performance relative to other models, its pre-training on massive web-scale corpora leaves it vulnerable to data contamination in widely used benchmarks (e.g., MATH-500). Consequently, experimental results derived from contaminated benchmarks on the Qwen2.5 series may be unreliable. To obtain trustworthy evaluation signals, we introduce a generator that creates fully synthetic arithmetic problems of arbitrary length and difficulty, yielding clean datasets we call RandomCalculation. Using these leakage-free datasets, we further show that under the RL protocol, only accurate reward signals yield steady improvements that surpass the model’s performance ceiling in mathematical reasoning, whereas noisy or incorrect rewards do not. Thus, we recommend that future studies evaluate on uncontaminated benchmarks and, when feasible, test various model series to ensure trustworthy conclusions about RL and related methods.
1 Introduction
In recent years, advances in reinforcement learning (RL) techniques have markedly strengthened the reasoning abilities of large language models (LLMs). Flagship systems, including OpenAI’s o1 (OpenAI, 2024b; 2025; a), DeepSeek-R1 (DeepSeek-AI et al., 2025), and QwQ (Qwen Team, 2024; 2025), now match or surpass human-level accuracy on a variety of challenging benchmarks. Among the contenders that excel on these mathematical benchmarks, the open-source Qwen family (Yang et al., 2024a; b; 2025), spanning parameters from 0.5B to 72B and maximally pre-trained on up to 36T high-quality tokens, yields state-of-the-art results in language understanding, mathematics, programming, and preference alignment.
Within this landscape, mathematical reasoning has emerged as a particularly discriminative test bed because it demands precise symbolic manipulation and multi-step logical deduction. Standard suites such as MATH-500 (Hendrycks et al., 2021), AIME (Li et al., 2024), AMC (Li et al., 2024) and Minerva Math (Lewkowycz et al., 2022) require models to parse natural-language problem statements, uncover the latent mathematical structure, and generate exact numeric answers. Recent work has further enhanced these capabilities by reinforcement learning with verifiable rewards (RLVR) (DeepSeek-AI et al., 2025): rule-based reward functions that return 1 when a predicted answer equals ground truth and 0 otherwise. Because the reward is computed analytically, RLVR removes the need for a separate learned reward model, lowering computational cost while providing an analytically precise training signal, especially attractive for domains like mathematics, where solutions are unambiguous.
Although RL nominally depends on accurate reward signals to guide effective training, recent studies (Zuo et al., 2025; Shao et al., 2025) report that even random or partially incorrect rewards can lift Qwen’s performance on standard math benchmarks, whereas the same procedures offer little or no benefit to Llama3.1-8B (Dubey et al., 2024) models. To understand why these seemingly reward-agnostic variants help Qwen but not Llama, we undertake a systematic comparison of the two model families under identical training protocols. We consider two working hypotheses to explain this phenomenon. (i) Data Contamination: Considering Qwen2.5 is pre-trained on massive web-scale corpora from the Internet, including GitHub repositories that store benchmark problems alongside their official solutions. If segments of the evaluation benchmarks leaked into Qwen’s pre-training corpus, spurious rewards could cue the model to retrieve memorized answers rather than acquire new reasoning skills. (ii) Strong Baseline Math Skills: Qwen’s pre-training endows it with higher mathematical capacity than Llama, so even noisy policy-gradient updates appear to help on MATH-500; however, if strong skills are the real driver, the same random-reward trick should still work on a clean benchmark. Distinguishing between these possibilities requires both a leakage audit and a rigorously out-of-distribution (OOD) RLVR evaluation, which we present in Section 4.2 and Section 4.3.
To assess the extent of potential data contamination in popular mathematical benchmarks, we evaluate two indicators: partial-prompt completion rate (can the model reconstruct the tail of a problem?) and partial-prompt answer accuracy (can the model give the correct answer with an incomplete problem?). As Figure 1 illustrates, we observe that Qwen can accurately complete the problem statement and provide the correct answer, whereas Llama does not. Specifically, prompting Qwen2.5-Math-7B with the first 60% of each MATH-500 question, we find that it regenerates the remaining 40% with a 54.60% exact-match rate and still answers 53.6% of these incomplete problems correctly. Llama3.1-8B, in contrast, scores 3.8% and 2.4% on both metrics. Crucially, on the newly released LiveMathBench (version 202505) (Liu et al., 2024), a post-hoc benchmark compiled after the release of the Qwen2.5 model family, Qwen’s completion rate drops sharply to 0.0%, consistent with Llama’s 0.0%. Its partial-prompt answer accuracy also falls to just 2.0%, comparable to Llama’s 1.0%. These results confirm that the earlier gains on MATH-500 may stem from memorized content rather than genuine reasoning. Hence, results derived from MATH-500 and similar datasets for Qwen models should be interpreted with caution.
Building on this evidence, we attribute that data contamination is the chief factor behind the seemingly “magical” success of random-reward or few-shot RLVR variants on Qwen models. To test this claim, we create a fully fresh benchmark (representative examples shown in Figure 2): We use an automatic generator to construct arithmetic expressions of arbitrary length with uniformly random operands and operators, guaranteeing that every instance post-dates the public release of Qwen2.5. A zero-shot evaluation on this benchmark shows no memorization: the accuracy of Qwen2.5 models declines monotonically with the number of computation steps, leaving ample room for improvement. To isolate the effect of reward quality, we next trained Qwen2.5-Math-7B under the standard RLVR protocol on two leakage-free subsets, respectively. The outcome is unambiguous: Correct rewards deliver consistent performance gains, surpassing the model’s performance ceiling; In contrast, random rewards make training highly unstable, yielding no reliable improvement, while inverse rewards rapidly erode the model’s mathematical-reasoning ability. These results rule out the “Strong Baseline Math Skills” explanation and directly implicate “Data Contamination”: once leakage is removed, the prior gains evaporate.
The contributions of our work can be summarized as follows:
We conduct a systematic leakage audit of widely used math benchmarks and demonstrate that Qwen’s sudden performance surge on MATH-500 under spurious-reward RLVR is chiefly attributable to unfair evaluation protocol, data contamination, and memorization rather than its strong math skills.
We design an automatic generator that creates arbitrarily long arithmetic expressions and a corresponding out-of-distribution evaluation protocol. Zero-shot results on this clean suite expose the absence of memorization and place Qwen and Llama on equal footing, enabling fair assessment of mathematical reasoning.
Using the leakage-free dataset, we conduct controlled RLVR experiments and demonstrate that only correct rewards yield stable performance gains, whereas random or inverse rewards provide no benefit, thereby highlighting the central role of reward fidelity. We recommend that future work rely on uncontaminated benchmarks or test various model series to draw trustworthy conclusions about RL-related methods.
2 Related Works
2.1 Reinforcement learning on Qwen2.5 for mathematical reasoning
A growing body of work investigates how reinforcement learning (RL) can amplify the mathematical-reasoning skills of the open-source Qwen2.5 family. Early studies used explicit verifiable rewards that score an answer as 1 / 0 by exact numerical agreement. Test-time RL (TTRL) (Zuo et al., 2025) applies this signal on-the-fly during inference and yields sizeable gains on MATH-500 and AIME2024. Subsequent efforts pursue extreme data efficiency: a single labeled example can suffice to boost performance (Wang et al., 2025a; see also the confidence-guided RLSC scheme of Li et al., 2025) or even to learn from an unlabeled instance via entropy minimization (Gao et al., 2025). A parallel line replaces external supervision with intrinsic signals derived from the model itself. Entropy-based objectives—RENT (Prabhudesai et al., 2025) and EM-RL (Agarwal et al., 2025)—reward low-entropy output distributions, while other methods such as Intuitor (Zhao et al., 2025b) and online self-training (Shafayat et al., 2025) rely on self-consistency or “self-certainty” as feedback. These approaches report large jumps on Qwen2.5-Math-7B, occasionally matching or surpassing stronger baselines (e.g. GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024)). Other variants explore noisy or partial rewards (Lv et al., 2025) and a label-free RLVR paradigm dubbed Absolute Zero (Zhao et al., 2025a). The generality of these gains, however, has been questioned. Zhang et al. (2025) observe an “underconfidence-to-overconfidence” trajectory in RLIF(Reinforcement Learning from Internal Feedback): accuracy rises early but later collapses as policy entropy shrinks. Shao et al. (2025) further shows that random or incorrect rewards can still improve Qwen2.5 yet fail to transfer to Llama or OLMo (OLMo et al., 2025), suggesting model-specific idiosyncrasies. Together, these findings indicate that while RL can unlock strong mathematical competence in Qwen2.5, its effectiveness hinges on careful reward design, entropy control, and underlying model capacity.
2.2 The Impact of Pretraining Corpora on the Reasoning Ability
The choice of pretraining corpora plays a crucial role in shaping the reasoning abilities of LLMs, particularly in mathematical domains. Several math-specific datasets (Paster et al., 2024; Han et al., 2024; Wang et al., 2024; Allal et al., 2025) have been proposed and have been shown to significantly enhance performance on relevant benchmarks. As highlighted in prior studies, models such as Qwen and Llama, despite having comparable size, exhibit notable differences in mathematical reasoning capabilities. For example, Qwen models tend to benefit more readily from reinforcement learning, whereas Llama models often show limited improvement under similar conditions. Wang et al. (2025b) found that mid-training Llama models on high-quality mathematical corpora substantially improve their performance, both at the base level and after reinforcement learning. These observations suggest that differences in pretraining data may be a key factor influencing a model’s receptiveness to reinforcement learning, and warrant further investigation.
3 Experimental Setup
3.1 Model Selection
Prior research on mathematical reasoning with large language models has focused predominantly on the Qwen-2.5 family (Yang et al., 2024a; b). Accordingly, we center our study on four representative checkpoints from this series: Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-Math-7B, and Qwen2.5-Math-7B-Instruct. For a controlled comparison, we also evaluate Llama3.1-8B and Llama3.1-8B-Instruct (Dubey et al., 2024), which possess comparable parameter counts and thus help isolate model-specific differences in behavior.
3.2 Memorization Capability
We provide the following details on evaluating the model’s memorization capability. Performance is assessed with two metrics: Partial-Prompt Completion Rate and Partial-Prompt Answer Accuracy.
3.2.1 Partial-Prompt Completion Rate
ROUGE Metric.
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) (Lin, 2004) is a family of metrics commonly used to evaluate automatic summarization and text generation systems. ROUGE score measures the overlap between the generated text and one or more reference texts, focusing primarily on recall. The most frequently used variants include:
ROUGE-N: Measures n-gram overlap. ROUGE-1 and ROUGE-2 refer to unigram and bigram recall respectively.
ROUGE-L: Measures the longest common subsequence (LCS) between the generated and reference text, capturing fluency and sentence-level structure.
ROUGE-Lsum: A variant of ROUGE-L adapted for multi-sentence summarization evaluation.
In this work, we utilize average ROUGE-L score to measure the model’s ability to reconstruct the remaining parts of a problem based on partial prefixes, serving as an indicator of the model’s memorization capacity.
Exact Match (EM).
EM is a binary accuracy metric that checks whether the model’s continuation exactly reproduces the reference. For each instance, let denote the model-generated continuation and denote the ground-truth continuation. if , ; otherwise . We first compute the rouge -L score between the and the , then assign
The final EM score is obtained by averaging over all test instances. Because ROUGE-L = 1 implies an exact, character-level match, higher EM values directly indicate a greater proportion of partial-prompts that the model recalls verbatim.
3.2.2 Partial-Prompt Answer Accuracy
Answer-Match Accuracy.
For each question, we supply the model with only a truncated prompt (e.g., the first of the original problem) and allow it to generate an unconstrained continuation. After generation, we check whether the completion contains the ground-truth answer; if so, the instance is scored as correct. Answer-Match Accuracy is defined as the fraction of prompts for which the model’s continuation embeds the correct answer. A high accuracy indicates that the model frequently “recovers” the answer even from a partial problem, which in turn may signal data contamination.
3.3 RLVR-Based Evaluation
The GRPO Algorithm.
We adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as our RLVR algorithm. Formally, for each question , GRPO samples a group of outputs from the old policy and then optimizes the policy model by maximizing the following objective:
where and are hyperparameters, and is the advantage calculated based on the relative rewards of the outputs inside each group only.
Reward Function for RandomCalculation.
The ground-truth answers to our randomly generated arithmetic problems often contain high-precision decimals. When the standard RLVR framework supplies only binary feedback (0 or 1), the model almost never receives positive reinforcement, making training highly unstable and prone to divergence. To address this limitation, we design a continuous reward that ranges from 0 to 1 and jointly penalizes both absolute and relative errors between the model prediction and the reference answer. This denser signal greatly stabilizes reinforcement learning. Let be the model output, be the reference answer, and be a small constant for numerical stability. The reward is computed as:
Reward Design.
Based on the configuration of Shao et al. (2025), we consider the following rewards for our RLVR experiments:
Correct: assigns 1 to a correct answer and 0 otherwise.
Random: assigns with probability and otherwise ( in our experiments).
Inverted: flips the correct signal, i.e., , so that correct solutions receive 0 and incorrect ones 1.
Mv-incorrect: retains only majority-voted incorrect labels and assigns 1 when the model output matches an incorrect label, and 0 otherwise.
4 Results & Analysis
4.1 Spurious Rewards on MATH-500
Following the work of Shao et al. (2025), we replicate the performance of Qwen2.5-Math-7B and Llama3.1-8B-Instruct on the MATH-500 benchmark under various reward signal configurations, using the same experimental setup. The training accuracy curves of MATH-500 are illustrated in Figure 3. Interestingly, the results demonstrate that while random rewards and mv-incorrect rewards noticeably boost accuracy for Qwen2.5-Math-7B, they have little or even adverse impacts on the performance of Llama3.1-8B-Instruct. Additionally, we apply the same RLVR procedure to Qwen2.5-Math-7B-Instruct and discover that the resulting gains are marginal when compared with those of Qwen2.5-Math-7B, indicating that the two Qwen variants exhibit differential sensitivity to RLVR.
Considering the base and instruct variants of Qwen are trained under different paradigms: the former is pre-trained as a general language model without exposure to any dialogue-specific templates, whereas the latter undergoes an additional instruction-tuning stage on data wrapped in a fixed dialogue template. This mismatch creates a training–testing gap for the base model at the start point of RLVR, and its initial accuracy is therefore likely underestimated. Consequently, to obtain a fair estimate of each model’s starting point, we next measure performance under four decoding configurations (w/ and w/o templates, with details shown in Table 1), and the corresponding results are summarized in Figure 4. Surprisingly, we discover that applying the official chat template substantially degrades performance for the Qwen base model: both Qwen-2.5-7B and Qwen-2.5-Math-7B suffer pronounced drops once the template is enabled, whether we use greedy or avg@16 sampling.
Building on this observation, we report two additional reference baselines in Figure 3: (i) a red dashed line, denoted Greedy (w/o Template), corresponds to the best performance of the initial model, and (ii) a grey dashed line, denoted pass@16 and adopted from Yue et al. (2025), serves as a plausible upper bound on the models’ initial performance. Viewed against these baselines, the seeming “RL gains” of Qwen-Math-7B largely reflect adaptation to the template format and merely converge to the Greedy (w/o Template) baseline, indicative of memory recall rather than genuine mathematical generalization. However, spurious rewards, e.g., random and mv-incorrect, still boost the accuracy of Qwen base and maintain the performance of Qwen instruct, while degrading Llama-Instruct eventually. This is the point we need to further explore in the following sections.
| Configuration | Do-Sample | Temperature | Top-P | Top-K | Chat Template |
| Greedy (w/o Template) | False | 1.0 | 1.0 | – | ✗ |
| Avg@16 (w/o template) | – | 0.7 | 0.8 | 20 | ✗ |
| Greedy (w/ Template) | False | 1.0 | 1.0 | – | ✓ |
| Avg@16 (w/ template) | – | 0.7 | 0.8 | 20 | ✓ |
Model Dataset Size 80%-Problem 60%-Problem 40%-Problem RougeL EM RougeL EM RougeL EM Qwen2.5-Math-7B MATH-500 500 81.25 65.80 78.06 54.60 69.01 39.20 AMC 83 77.38 55.42 70.25 42.17 75.17 36.14 AIME2024 30 74.04 56.67 55.31 20.00 57.72 16.67 AIME2025 30 54.71 16.67 34.88 0.00 27.43 0.00 MinervaMath 272 36.08 2.94 31.22 0.37 29.35 0.00 LiveMathBench 100 42.76 5.00 32.78 0.00 29.97 0.00 Qwen2.5-7B MATH-500 500 66.42 40.20 60.98 21.20 50.36 8.20 AMC 83 73.24 49.40 64.42 33.73 63.79 28.92 AIME2024 30 59.80 30.00 48.69 13.33 44.65 10.00 AIME2025 30 54.61 10.00 37.59 0.00 30.30 0.00 MinervaMath 272 35.24 2.94 32.35 0.37 27.89 0.00 LiveMathBench 100 41.15 4.00 32.74 0.00 27.95 0.00 Llama3.1-8B MATH-500 500 48.33 17.80 40.55 3.80 32.07 0.60 AMC 83 44.54 4.82 30.62 0.00 27.10 0.00 AIME2024 30 50.50 13.33 30.80 0.00 26.08 0.00 AIME2025 30 47.04 10.00 33.49 0.00 25.20 0.00 MinervaMath 272 36.24 2.21 29.52 0.00 27.11 0.00 LiveMathBench 100 35.55 5.00 31.93 0.00 26.88 0.00
4.2 Analysis of Memorization Capability Results
Considering the Qwen series is trained on massive web-scale corpora, we hypothesize that its divergent RLVR behavior from Llama is because the evaluation set MATH-500 may be inadvertently contaminated in Qwen’s large-scale training data, which is hard to eliminate completely. To verify our hypothesis, we probe memorization on several widely used mathematical-reasoning benchmarks. Concretely, we truncate the original questions at 40%, 60%, and 80% of their lengths and feed these partial questions as prompts into the model, and then evaluate the model’s partial-prompt completion rate by computing ROUGE and EM scores (details in 3.2) between the generated completion and ground-truth continuations. In addition, we evaluate the model’s partial-prompt answer accuracy by checking if the continuation contains the correct answer, across both partial and full question settings.
The detailed results are presented in Table 2, revealing strong signs of data contamination in the Qwen2.5 series models when evaluated on commonly used benchmarks, such as MATH-500, AMC, and AIME2024. For instance, when only the first 60% of the questions are provided, Qwen2.5-Math-7B is able to accurately reconstruct more than half of the remaining problems on MATH-500 with greedy decoding, achieving a completion rate of 54.6%. Even when just 40% proportion of the questions are shown, the model still manages to recover 39.2% of the unseen problems on MATH-500. Similar patterns are observed on AMC and AIME2024. These results indicate that the evaluation benchmarks for Qwen2.5 may suffer from data contamination: Although pre-training on massive web-scale corpora brings strong capacity on mathematical reasoning, e.g., superior performance on recently introduced mathematical tasks like LiveMathBench and AIME2025, those large-scale corpora also include publicly available benchmark problems inevitably, leading to less convincing results of old benchmarks, while removing such instances during large-scale crawling is notoriously difficult.
Meanwhile, we summarize the answer accuracy of the model’s output under different proportions of input questions. The detailed results are shown in Figure 5. We observe that under greedy decoding, even when partial questions are provided, the Qwen2.5 models achieve remarkably high accuracy on MATH-500. For instance, with 80% proportion of the questions, Qwen2.5-Math-7B reaches an accuracy of 63.8% on MATH-500. Even when only 40% proportion of the questions are given, the model still achieves 41.2% accuracy. Because our evaluation matches only the final numeric answer, the model can sometimes “accidentally” output the correct value despite faulty reasoning, which explains the anomalous jump observed for Llama on AIME2025 when only 40% of the questions are provided. Among the several questions that the model solves correctly, the responses contain coherent reasoning chains and even syntactically valid Python code, which, however, is not executed. And several examples are shown in Figure 8 to 12. The emergence of such structured solutions (also observed in Shao et al. (2025)) indicates that the training corpora may have included publicly available resources where benchmark problems are accompanied by detailed solutions. Therefore, the sudden performance surge under random rewards or even mv-incorrect rewards, observed in Figure 3, arises because such rewards inadvertently trigger Qwen to retrieve its memorized answers from the pre-training corpus owing to the clipping bias of GRPO (Shao et al., 2025), rather than stimulating Qwen’s existing distribution of reasoning patterns as explained in Shao et al. (2025).
4.3 Spurious Rewards on RandomCalculation
To further support our hypothesis that the anomalous performance surge of the Qwen2.5-Math-7B on the MATH-500 benchmark is primarily caused by data contamination rather than the model’s intrinsic mathematical reasoning ability, we replicate this experiment on a newly constructed dataset that the model has never encountered before. We hypothesize that for math problems free from contamination, the model’s reasoning ability still requires properly aligned reward signals to yield meaningful performance improvements.
RandomCalculation Dataset Construction.
To obtain an uncontaminated evaluation benchmark, we employ Algorithm 1 to construct a suite of mathematically challenging yet verifiable computation datasets. These datasets are composed of expressions built from basic numerical elements, including integers from 0 to 100, as well as fractions, squares, and cubes derived from them. Using these components, we randomly generate mathematical expressions involving between 1 and 20 steps, utilizing the four fundamental arithmetic operations: addition, subtraction, multiplication, and division. To construct the final datasets, we append a standardized problem prefix to each generated expression, resulting in 20 sub-datasets, each containing 1,000 unique problems. We refer to this suite of datasets as RandomCalculation. Examples from the datasets can be found in Figure 2.
Qwen2.5’s initial performance on RandomCalculation.
We first test the zero-shot performance of the Qwen2.5 series models on the RandomCalculation dataset, with the results shown in Figure 6. We find that when using chat templates, the models’ mathematical reasoning ability tends to degrade as the number of computation steps increases, leaving ample room for improvement in multi-step calculation problems. When chat templates are removed, the reasoning performance peaks on problems involving three computation steps, and then gradually declines.
RLVR results on RandomCalculation.
We conduct RLVR training on Qwen2.5-Math-7B using the RandomCalculation datasets, aiming to guide the model toward producing answers that are as close as possible to the correct ones. Experiments are set on two datasets comprising 5-step and 10-step calculation problems, respectively. Each dataset contains 1,000 problems, with 700 used for training and the remaining 300 reserved for validation. The training progress is shown in Figure 7, showing that under correct rewards, the model’s performance steadily improves throughout training. With random or incorrect rewards, training becomes unstable and inconsistent. Under inverted rewards, the model collapses rapidly. These findings suggest that for problems not leaked during pretraining, only correct reward signals can effectively guide the model toward improved reasoning performance. For comparison, we also evaluate Llama3.1-8B-Instruct and observe consistent findings: Only accurate reward signals consistently yield stable performance gains.
Qwen’s stronger mathematical capabilities.
Considering obtaining a reward of 1 on a RandomCalculation instance is virtually impossible, we report Max@16, the highest score among 16 samples, instead of Pass@16 in Figure 7. We observe that on our RandomCalculation datasets, Qwen2.5-Math-7B can surpass the Max@16 threshold when provided with correct reward signals. This finding indicates that reward-aligned RLVR effectively transfers the model’s high-accuracy single-step arithmetic skills (see template-enabled accuracy in Figure 6) to more complex multi-step calculations, as illustrated by one reasoning trace in Figure 13. In contrast, under the incorrect and random reward configurations, the Qwen model either maintains its base performance or exhibits only marginal and unstable improvements. This further highlights the critical role of correct and well-aligned reward signals in enhancing model performance on uncontaminated datasets. Notably, Llama3.1-8B-Instruct fails to surpass the Max@16 threshold even when trained with correct reward signals, and its accuracy falls below the greedy-decoding baseline when exposed to spurious signals. This discrepancy indirectly suggests that Qwen2.5 exhibits stronger mathematical capabilities compared to Llama3.1; however, such inherent strength is not the root cause of its performance boosting under spurious reward on contaminated datasets.
5 Discussion
Due to limited computational resources, our experiments were restricted to a subset of commonly used Qwen2.5 series models. The Qwen3 series will be included in our future experimental evaluation. As the successor to Qwen2.5, it is likely that Qwen3 still suffers from similar issues. Many reinforcement learning algorithms have recently been proposed to enhance mathematical reasoning capabilities. Given the rapid pace of development, it is infeasible to conduct a comprehensive evaluation of all these methods in the short term. In future work, our efforts will focus on systematically evaluating a more diverse set of benchmarks, reinforcement learning methods, and model families. In parallel, we will carry out an in-depth theoretical investigation of this unexpected behavior, with the goal of shedding light on its root causes and advancing our understanding of the underlying mechanisms.
6 Conclusion
Recent studies have shown that the mathematical reasoning performance of Qwen2.5 models can be substantially improved with minimal supervision, ranging from few-shot prompts to single examples or even without additional data. Some studies further suggest that the mathematical reasoning abilities of the Qwen2.5 model can be improved with reward signals that are random or incorrect. While the effectiveness of these methods is striking, their gains appear largely limited to the Qwen2.5 series. This distinctive phenomenon motivated our study, which reveals that the large-scale pretraining corpus of the Qwen2.5 series models may be contaminated by data that overlaps with several widely used mathematical benchmarks, including MATH-500, AMC, and AIME 2024. We suspect that such data contamination is a key factor for Qwen’s improvement on MATH-500 under those spurious rewards. Moreover, to ensure reliable evaluation of RL-based algorithms, we construct a leakage-free RandomCalculation dataset composed of expressions built from basic numerical elements, and a zero-shot test on this dataset shows Qwen has ample room for improvement in multi-step calculation problems. Leveraging this contamination-free dataset, we run carefully controlled RLVR experiments and find that only accurate reward signals deliver consistent performance improvements, while noisy or reversed rewards yield no measurable gains, underscoring the pivotal importance of reward integrity. We therefore recommend that future research conduct evaluations on clean benchmarks and assess multiple model families to ensure their RL-based methods are robust.
Acknowledgments
Qin Liu is supported by the Amazon Nova Trusted AI Prize.
Appendix A Appendix
| Model | Configuration | 100% | 80% | 60% | 40% |
| Qwen2.5-Math-7B | Greedy (w/o Template) | 72.20 | 63.80 | 53.60 | 41.20 |
| Avg@16 (w/o Template) | 68.53 | 57.25 | 45.51 | 31.03 | |
| Greedy (w/ Template) | 50.60 | 29.20 | 20.20 | 10.00 | |
| Avg@16 (w/ Template) | 48.84 | 28.01 | 18.99 | 10.36 | |
| Qwen2.5-Math-7B-Instruct | Greedy (w/o Template) | 83.60 | 53.80 | 31.80 | 16.00 |
| Avg@16 (w/o Template) | 81.76 | 50.80 | 31.54 | 15.43 | |
| Greedy (w/ Template) | 82.20 | 3.80 | 24.40 | 13.20 | |
| Avg@16 (w/ Template) | 82.01 | 43.66 | 25.14 | 12.78 | |
| Qwen2.5-7B | Greedy (w/o Template) | 67.60 | 49.60 | 30.80 | 18.40 |
| Avg@16 (w/o Template) | 66.20 | 46.04 | 30.01 | 16.10 | |
| Greedy (w/ Template) | 40.00 | 22.20 | 13.40 | 6.80 | |
| Avg@16 (w/ Template) | 38.15 | 21.75 | 11.89 | 5.25 | |
| Qwen2.5-7B-Instruct | Greedy (w/o Template) | 72.80 | 50.00 | 31.20 | 15.20 |
| Avg@16 (w/o Template) | 74.90 | 50.18 | 30.88 | 15.79 | |
| Greedy (w/ Template) | 72.20 | 36.00 | 20.80 | 10.00 | |
| Avg@16 (w/ Template) | 73.69 | 37.16 | 20.67 | 8.75 | |
| Llama3.1-8B | Greedy (w/o Template) | 1.60 | 2.00 | 2.40 | 2.00 |
| Avg@16 (w/o Template) | 3.12 | 2.81 | 2.80 | 2.30 | |
| Greedy (w/ Template) | – | – | – | – | |
| Avg@16 (w/ Template) | – | – | – | – | |
| Llama3.1-8B-Instruct | Greedy (w/o Template) | 43.60 | 24.80 | 15.00 | 7.40 |
| Avg@16 (w/o Template) | 39.61 | 24.76 | 14.21 | 6.84 | |
| Greedy (w/ Template) | 38.80 | 17.20 | 10.60 | 5.00 | |
| Avg@16 (w/ Template) | 38.24 | 18.46 | 9.95 | 4.25 |
| Model | Dataset | Size | 100% | 80% | 60% | 40% |
| Qwen2.5-Math-7B | MATH-500 | 500 | 72.20 | 63.80 | 53.60 | 41.20 |
| AMC | 83 | 53.01 | 53.01 | 43.37 | 43.37 | |
| AIME2024 | 30 | 13.33 | 13.33 | 13.33 | 13.33 | |
| AIME2025 | 30 | 16.67 | 3.33 | 0.00 | 0.00 | |
| MinervaMath | 272 | 14.71 | 6.25 | 2.94 | 2.57 | |
| LiveMathBench | 100 | 7.00 | 9.00 | 2.00 | 3.00 | |
| Qwen2.5-Math-7B-Instruct | MATH-500 | 500 | 83.60 | 53.80 | 31.80 | 16.00 |
| AMC | 83 | 57.83 | 27.71 | 7.23 | 4.82 | |
| AIME2024 | 30 | 16.67 | 3.33 | 3.33 | 0.00 | |
| AIME2025 | 30 | 6.67 | 0.00 | 0.00 | 0.00 | |
| MinervaMath | 272 | 20.96 | 7.72 | 3.68 | 1.84 | |
| LiveMathBench | 100 | 9.00 | 4.00 | 4.00 | 1.00 | |
| Qwen2.5-7B | MATH-500 | 500 | 67.60 | 49.60 | 30.80 | 18.40 |
| AMC | 83 | 43.37 | 32.53 | 24.10 | 22.89 | |
| AIME2024 | 30 | 6.67 | 6.67 | 6.67 | 3.33 | |
| AIME2025 | 30 | 3.33 | 0.00 | 0.00 | 0.00 | |
| MinervaMath | 272 | 12.87 | 4.41 | 2.57 | 2.21 | |
| LiveMathBench | 100 | 9.00 | 8.00 | 3.00 | 4.00 | |
| Qwen2.5-7B-Instruct | MATH-500 | 500 | 72.80 | 50.00 | 31.20 | 15.20 |
| AMC | 83 | 44.58 | 34.94 | 22.89 | 14.46 | |
| AIME2024 | 30 | 16.67 | 6.67 | 0.00 | 0.00 | |
| AIME2025 | 30 | 6.67 | 3.33 | 0.00 | 0.00 | |
| MinervaMath | 272 | 18.01 | 5.51 | 3.68 | 2.21 | |
| LiveMathBench | 100 | 9.00 | 6.00 | 2.00 | 4.00 | |
| Llama3.1-8B | MATH-500 | 500 | 1.60 | 2.00 | 2.40 | 2.00 |
| AMC | 83 | 4.82 | 2.41 | 1.20 | 2.41 | |
| AIME2024 | 30 | 0.00 | 0.00 | 0.00 | 0.00 | |
| AIME2025 | 30 | 0.00 | 0.00 | 0.00 | 3.33 | |
| MinervaMath | 272 | 2.57 | 1.10 | 1.10 | 0.00 | |
| LiveMathBench | 100 | 0.00 | 2.00 | 1.00 | 1.00 | |
| Llama3.1-8B-Instruct | MATH-500 | 500 | 43.60 | 24.80 | 15.00 | 7.40 |
| AMC | 83 | 21.69 | 10.84 | 3.61 | 2.41 | |
| AIME2024 | 30 | 10.00 | 0.00 | 0.00 | 0.00 | |
| AIME2025 | 30 | 0.00 | 0.00 | 0.00 | 0.00 | |
| MinervaMath | 272 | 10.29 | 4.04 | 2.57 | 1.84 | |
| LiveMathBench | 100 | 3.00 | 3.00 | 2.00 | 0.00 |
| Model | Dataset | Size | 100% | 80% | 60% | 40% |
| Qwen2.5-Math-7B | MATH-500 | 500 | 50.60 | 29.20 | 20.20 | 10.00 |
| AMC | 83 | 37.35 | 20.48 | 14.46 | 7.23 | |
| AIME2024 | 30 | 10.00 | 6.67 | 0.00 | 0.00 | |
| AIME2025 | 30 | 6.67 | 3.33 | 0.00 | 0.00 | |
| MinervaMath | 272 | 9.19 | 4.78 | 2.94 | 2.21 | |
| LiveMathBench | 100 | 5.00 | 8.00 | 2.00 | 2.00 | |
| Qwen2.5-Math-7B-Instruct | MATH-500 | 500 | 82.20 | 43.80 | 24.40 | 13.20 |
| AMC | 83 | 55.42 | 14.46 | 7.23 | 4.82 | |
| AIME2024 | 30 | 20.00 | 0.00 | 3.33 | 0.00 | |
| AIME2025 | 30 | 16.67 | 0.00 | 0.00 | 0.00 | |
| MinervaMath | 272 | 26.47 | 6.62 | 4.78 | 1.84 | |
| LiveMathBench | 100 | 8.00 | 12.00 | 6.00 | 5.00 | |
| Qwen2.5-7B | MATH-500 | 500 | 40.00 | 22.20 | 13.40 | 6.80 |
| AMC | 83 | 27.71 | 9.64 | 3.61 | 3.61 | |
| AIME2024 | 30 | 6.67 | 0.00 | 0.00 | 0.00 | |
| AIME2025 | 30 | 6.67 | 0.00 | 0.00 | 0.00 | |
| MinervaMath | 272 | 8.09 | 4.41 | 2.21 | 1.47 | |
| LiveMathBench | 100 | 8.00 | 6.00 | 2.00 | 1.00 | |
| Qwen2.5-7B-Instruct | MATH-500 | 500 | 72.20 | 36.00 | 20.80 | 10.00 |
| AMC | 83 | 48.19 | 10.84 | 4.82 | 2.41 | |
| AIME2024 | 30 | 6.67 | 3.33 | 0.00 | 0.00 | |
| AIME2025 | 30 | 6.67 | 0.00 | 0.00 | 0.00 | |
| MinervaMath | 272 | 23.53 | 6.25 | 3.68 | 2.94 | |
| LiveMathBench | 100 | 10.00 | 10.00 | 3.00 | 2.00 | |
| Llama3.1-8B | MATH-500 | 500 | – | – | – | – |
| AMC | 83 | – | – | – | – | |
| AIME2024 | 30 | – | – | – | – | |
| AIME2025 | 30 | – | – | – | – | |
| MinervaMath | 272 | – | – | – | – | |
| LiveMathBench | 100 | – | – | – | – | |
| Llama3.1-8B-Instruct | MATH-500 | 500 | 38.80 | 17.20 | 10.60 | 5.00 |
| AMC | 83 | 25.30 | 6.02 | 2.41 | 1.20 | |
| AIME2024 | 30 | 6.67 | 0.00 | 0.00 | 0.00 | |
| AIME2025 | 30 | 0.00 | 0.00 | 3.33 | 0.00 | |
| MinervaMath | 272 | 15.81 | 2.94 | 3.68 | 2.21 | |
| LiveMathBench | 100 | 2.00 | 2.00 | 4.00 | 1.00 |
| Model | Dataset | Size | 100% | 80% | 60% | 40% |
| Qwen2.5-Math-7B | MATH-500 | 500 | 68.53 | 57.25 | 45.51 | 31.03 |
| AMC | 83 | 49.47 | 48.57 | 40.06 | 37.27 | |
| AIME2024 | 30 | 18.33 | 13.75 | 14.58 | 14.17 | |
| AIME2025 | 30 | 6.88 | 1.67 | 0.83 | 0.00 | |
| MinervaMath | 272 | 11.40 | 5.12 | 3.29 | 1.77 | |
| LiveMathBench | 100 | 9.50 | 7.56 | 3.38 | 3.06 | |
| Qwen2.5-Math-7B-Instruct | MATH-500 | 500 | 81.76 | 50.80 | 31.54 | 15.43 |
| AMC | 83 | 51.28 | 28.54 | 9.41 | 4.52 | |
| AIME2024 | 30 | 12.29 | 4.58 | 1.88 | 0.00 | |
| AIME2025 | 30 | 10.42 | 1.46 | 1.04 | 0.42 | |
| MinervaMath | 272 | 18.50 | 7.08 | 4.07 | 1.75 | |
| LiveMathBench | 100 | 10.88 | 7.81 | 5.38 | 2.56 | |
| Qwen2.5-7B | MATH-500 | 500 | 66.20 | 46.04 | 30.01 | 16.10 |
| AMC | 83 | 40.06 | 32.23 | 27.71 | 22.52 | |
| AIME2024 | 30 | 11.04 | 7.50 | 4.58 | 4.58 | |
| AIME2025 | 30 | 7.92 | 0.83 | 0.42 | 0.21 | |
| MinervaMath | 272 | 10.66 | 5.51 | 3.26 | 1.38 | |
| LiveMathBench | 100 | 7.75 | 6.50 | 4.06 | 3.19 | |
| Qwen2.5-7B-Instruct | MATH-500 | 500 | 74.90 | 50.18 | 30.88 | 15.79 |
| AMC | 83 | 43.07 | 33.51 | 17.62 | 11.75 | |
| AIME2024 | 30 | 11.67 | 5.42 | 1.04 | 0.21 | |
| AIME2025 | 30 | 6.67 | 1.04 | 0.42 | 0.00 | |
| MinervaMath | 272 | 18.18 | 6.82 | 4.11 | 1.59 | |
| LiveMathBench | 100 | 10.06 | 7.75 | 5.12 | 3.31 | |
| Llama3.1-8B | MATH-500 | 500 | 3.12 | 2.81 | 2.80 | 2.30 |
| AMC | 83 | 0.98 | 1.58 | 1.66 | 1.88 | |
| AIME2024 | 30 | 0.00 | 0.00 | 0.00 | 0.21 | |
| AIME2025 | 30 | 0.00 | 0.00 | 0.21 | 0.21 | |
| MinervaMath | 272 | 1.93 | 1.24 | 0.92 | 1.06 | |
| LiveMathBench | 100 | 0.25 | 1.38 | 0.44 | 1.06 | |
| Llama3.1-8B-Instruct | MATH-500 | 500 | 39.61 | 24.76 | 14.21 | 6.84 |
| AMC | 83 | 21.84 | 9.26 | 3.31 | 2.48 | |
| AIME2024 | 30 | 5.42 | 1.46 | 0.42 | 0.21 | |
| AIME2025 | 30 | 0.42 | 0.42 | 0.00 | 0.00 | |
| MinervaMath | 272 | 8.89 | 5.08 | 3.19 | 1.45 | |
| LiveMathBench | 100 | 3.25 | 2.00 | 1.38 | 1.88 |
| Model | Dataset | Size | 100% | 80% | 60% | 40% |
| Qwen2.5-Math-7B | MATH-500 | 500 | 48.84 | 28.01 | 18.99 | 10.36 |
| AMC | 83 | 35.32 | 14.83 | 10.99 | 8.36 | |
| AIME2024 | 30 | 11.46 | 5.83 | 0.83 | 1.25 | |
| AIME2025 | 30 | 3.54 | 1.46 | 0.42 | 0.00 | |
| MinervaMath | 272 | 8.94 | 4.83 | 2.96 | 1.70 | |
| LiveMathBench | 100 | 5.75 | 4.56 | 3.06 | 2.06 | |
| Qwen2.5-Math-7B-Instruct | MATH-500 | 500 | 82.01 | 43.66 | 25.14 | 12.78 |
| AMC | 83 | 50.53 | 14.68 | 6.70 | 4.07 | |
| AIME2024 | 30 | 12.50 | 0.83 | 0.21 | 0.00 | |
| AIME2025 | 30 | 10.62 | 0.00 | 0.42 | 0.00 | |
| MinervaMath | 272 | 27.07 | 7.01 | 4.76 | 1.93 | |
| LiveMathBench | 100 | 11.19 | 10.94 | 6.56 | 4.44 | |
| Qwen2.5-7B | MATH-500 | 500 | 38.15 | 21.75 | 11.89 | 5.25 |
| AMC | 83 | 20.78 | 8.89 | 4.22 | 2.86 | |
| AIME2024 | 30 | 5.42 | 0.00 | 0.00 | 0.21 | |
| AIME2025 | 30 | 0.62 | 0.21 | 0.00 | 0.00 | |
| MinervaMath | 272 | 8.18 | 3.52 | 2.27 | 1.77 | |
| LiveMathBench | 100 | 5.00 | 3.69 | 2.81 | 1.12 | |
| Qwen2.5-7B-Instruct | MATH-500 | 500 | 73.69 | 37.16 | 20.67 | 8.75 |
| AMC | 83 | 44.80 | 10.54 | 5.12 | 1.88 | |
| AIME2024 | 30 | 12.92 | 1.25 | 1.25 | 0.00 | |
| AIME2025 | 30 | 5.21 | 0.00 | 0.00 | 0.00 | |
| MinervaMath | 272 | 24.95 | 6.20 | 3.56 | 1.79 | |
| LiveMathBench | 100 | 10.25 | 8.75 | 4.31 | 3.56 | |
| Llama3.1-8B | MATH-500 | 500 | – | – | – | – |
| AMC | 83 | – | – | – | – | |
| AIME2024 | 30 | – | – | – | – | |
| AIME2025 | 30 | – | – | – | – | |
| MinervaMath | 272 | – | – | – | – | |
| LiveMathBench | 100 | – | – | – | – | |
| Llama3.1-8B-Instruct | MATH-500 | 500 | 38.24 | 18.46 | 9.95 | 4.25 |
| AMC | 83 | 19.43 | 6.10 | 3.24 | 2.48 | |
| AIME2024 | 30 | 6.25 | 0.42 | 0.00 | 0.00 | |
| AIME2025 | 30 | 0.42 | 0.00 | 0.21 | 0.21 | |
| MinervaMath | 272 | 14.38 | 3.15 | 2.44 | 1.47 | |
| LiveMathBench | 100 | 3.00 | 2.56 | 1.69 | 2.06 |